OpenStudy (anonymous):

5-(4/k)-(1/k^2)=0 show work

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 5-\frac{4}{k} -\frac{1}{k^2}=0 }\)

OpenStudy (solomonzelman):

You can multiply each times \(k^2\), and you will get a quadratic equation.

OpenStudy (solomonzelman):

I meant each *term* times \(k^2\)

OpenStudy (anonymous):

can you show me

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 5\color{red}{\times k^2}-\frac{4}{k} \color{red}{\times k^2}-\frac{1}{k^2}\color{red}{\times k^2}=0 }\)

OpenStudy (solomonzelman):

then simplify...

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 5 k^2-4k-1=0 }\)

OpenStudy (anonymous):

is that the answer you get?

OpenStudy (solomonzelman):

That is not the answer, all I have done is rearranged/worked your equation into a quadratic equation.

OpenStudy (solomonzelman):

Now you have to solve it. You can apply the *quadratic formula*

OpenStudy (anonymous):

oh ok